OpenStudy (anonymous):

Rational equation problem: A number minus 5 times the reciprocal of the number is 4. Find the number

OpenStudy (mathmale):

Let the number be x. 1. What is the reciprocal of the number? Express it algebraically. 2. What is "5 times the reciprocal of the number?" Express it algebraically. 3. what is "the number less 5 rimes the reciprocal of the number?"

OpenStudy (mrnood):

This could be interpreted 2 ways: (A number minus 5) times (the reciprocal of the number) is 4. OR (A number) minus (5 times the reciprocal of the number) is 4. No real way of telling from the question which is intended...

OpenStudy (mathmale):

Bonus to anyone who attempts to solve both possible interpretations of this problem AND shows all work. Lily: Mind answering my questions (above)?

OpenStudy (mrnood):

(A number minus 5) times (the reciprocal of the number) is 4. ( x-5 ) * ( 1/x ) = 4 (A number) minus (5 times the reciprocal of the number) is 4. ( x ) - ( 5 * 1/x )= 4